# A Comprehensive Survey on World Models for Embodied AI

### v1

@@ 3.1.1 Sequential Simulation and Inference
Ha and Schmidhuber [9] introduced an early world model that encodes observations into a latent space and employs an RNN to model dynamics for policy optimization. Building on this, PlaNet [38] introduced the Recurrent State-Space Model (RSSM), which fuses deterministic memory with stochastic components to enable robust long-horizon imagination. The successor models Dreamer, DreamerV2, and DreamerV3 [10, 11, 12] further advanced this formulation, inspiring a broad line of subsequent research.
@@
Building on RSSM, several variants modify or eliminate the decoder to better capture dynamics. For example, Dreaming [110] uses contrastive learning and linear methods to mitigate state shifts, whereas DreamerPro [111] replaces the decoder with prototypes to suppress visual distractions. To further enhance robustness, HRSSM [25] was proposed, featuring a dual-branch architecture that aligns latent observations and shares information without reconstruction. Beyond architectural refinements, DisWM [112] disentangles semantic knowledge from video content, distilling it into a world model that enables cross-domain generalization.
@@
A unifying theme across recent RSSM extensions is transferability, which reflects generalization across modalities, tasks, and embodiments for robust real-world robotics. At the representation level, PreLAR [52] learns implicit action abstractions to bridge video-pretrained representations and control fine-tuning. Similarly, Wang et al. [113] used optical flow as an embodiment-agnostic action representation to refine behavior-cloned policies, facilitating transfer across diverse embodiments. SENSEI [114] distilled a Vision-Language Model (VLM) to derive semantic rewards and employed an RSSM that learns to predict and propagate these rewards internally. Under limited supervision, SR-AIF [115] exploits prior preference learning and self-revision to enable adaptive learning in sparse-reward, continuous-control settings. To mitigate the Sim-to-Real (S2R) gap, ReDRAW [116] is pretrained in simulation and adapted to the real environment using a limited amount of reward-free data, applying residual corrections to the latent dynamics. To handle mismatches, AdaWM [117] identifies discrepancies between the learned dynamics and the planner and selectively fine-tunes critical components. Other methods like WMP [118] address S2R transfer for challenging tasks, and DayDreamer [43] demonstrated sample-efficient deployment on physical robots. To broaden transfer, FOUNDER [119] grounds representations from foundation models in the world-model state space, using temporal-distance prediction to handle flexible goals, and LUMOS [120] introduced a language-conditioned imitation framework that operates on-policy in latent space with intrinsic rewards, enabling zero-shot transfer to real-world robotics.
@@
RSSM-based models have also been developed for autonomous driving. MILE [81] leverages offline expert data to enable imagined future states for planning. SEM2 [83] integrates semantic filtering with multi-source sampling to extract driving-relevant features and balance data distribution. Popov et al. [121] addressed covariate shift through a latent generative world model that realigns policies with expert states. For safety, VL-SAFE [122] supervises world models using safety scores derived from a VLM to generate safe trajectories. Finally, CALL [123] extended the RSSM framework to Multi-Agent RL by introducing ego-centric information sharing to enhance planning capabilities.
@@
In contrast to RSSM, TransDreamer [28] introduced a Transformer State-Space Model (TSSM) that replaces the recurrent core in Dreamer, thereby substantially enhancing its capacity to capture long-horizon dependencies. The complementary OSVI-WM [124] employs a causal Transformer for one-shot Imitation Learning (IL), autoregressively predicting the future latent trajectory and decoding it into physical waypoints for robot control.
@@
Some approaches continue to employ RNNs to capture temporal dependencies. On the modeling side, RWM [125] introduced a dual-autoregressive, domain-agnostic neural simulator for long-horizon prediction. X-MOBILITY [126], in contrast, disentangles modeling from policy learning, using multi-head decoders for large-scale pretraining followed by supervised fine-tuning to derive strong policies. For humanoid locomotion, DWL [48] and WMR [127] adopted end-to-end (E2E) frameworks. These frameworks reconstruct states from partial observations using either denoising or a gradient-blocked state estimator, which enables zero-shot transfer across complex real-world terrains.
@@
Recently, State Space Models (SSMs), exemplified by Mamba, have emerged as alternatives to RNNs and Transformers, combining linear-time complexity with long-horizon modeling capacity. Building on this, GLAM [57] improves both fidelity and efficiency via a Mamba-based parallel framework that integrates global and local modules to capture contextual and fine-grained dynamics.
@@
Beyond forward temporal modeling, Inverse Dynamics Modeling (IDM) is a key paradigm in world model construction. An IDM infers the actions required to transition between an initial and a target state. Agrawal et al. [128] integrated a forward model with an IDM for multi-step prediction, establishing the basis for subsequent research. More recent work includes GLAMOR [39], which trains an object-conditioned IDM to predict the actions necessary to reach a specified target. In Dreamer-style agents, Iso-Dream[40] leverages IDM to decompose the world model into controllable and uncontrollable components, using the rollout of uncontrollable states to guide policy learning.
@@
Recent RSSM-centric studies have begun to exploit token-level dependencies to strengthen representation learning and temporal reasoning. For example, MWM [41] decouples visual tokens from RSSM-based dynamics via a masked autoencoder, improving both performance and data efficiency. NavMorph [129] introduced a self-evolving RSSM with a contextual evolution memory for online adaptation. For temporal abstraction, WISTER [130] employed action-conditioned contrastive predictive coding to train a TSSM that captures high-level temporal features. Similarly, TWM [29] used a Transformer to align multimodal tokens with historical states during training, while relying on a lightweight policy at inference. To handle long-horizon tasks, some approaches integrate LLMs with RSSMs to decompose objectives into subtasks. EvoAgent [131], for example, uses an LLM to guide low-level actions and regularizes RSSM updates. RoboHorizon [132], in contrast, enhances task recognition with dense rewards and leverages key task segments via a masked autoencoder.
@@
In autonomous driving, token-based sequence representations are increasingly adopted to model cross-modal interactions and spatiotemporal structures. DrivingWorld [133] pairs next-state prediction for temporal dynamics with next-token prediction for spatial structure. For multimodal control, Doe-1 [134] formulates closed-loop driving as autoregressive prediction over perception-description-action tokens, which unifies perception, prediction, and planning, and DrivingGPT [23] interleaves vision and action tokens and casts world modeling and trajectory planning as next-token prediction. To enhance diversity and safety, LatentDriver [135] models future actions as a mixture distribution and actuates the world model with planner-sampled intermediate actions. At the same time, Vasudevan et al. [136] proposed an adaptive model that predicts surrounding agents for safe navigation.
@@
The token-based paradigm also extends to broader robotics. Within RL, IRIS [44] and TWM [137] leverage discrete tokens to enable data-efficient policy learning via imagined or hybrid rollouts. DyWA [138] improves action learning by conditioning on trajectory dynamics and jointly predicting future states with single-view point-cloud and proprioceptive modalities. EgoAgent [139] interleaves state-action sequence modeling within a Transformer, enabling unified perception, prediction, and action inference. Tokenized representations unify multimodal inputs, including vision, language, and action (VLA), enabling generalist agents with cross-domain adaptability, as shown in WorldVLA [67]. Recent studies encode environmental states as discrete symbolic tokens and condition next-token prediction on action, as demonstrated by DCWM [140] and TrajWorld [141].
@@
Recent studies have strengthened the link between tokenized representations and planning, particularly through object-centric approaches. These models, such as CarFormer [142], the work of Jeong et al. [143], and Dyn-O [69], represent scenes as a collection of slots. CarFormer autoregressively models the relationships between these slots in BEV. Jeong et al. added language-guided manipulation, while Dyn-O used a Mamba with dropout scheduling for temporal modeling and to disentangle static from dynamic elements. $\Delta$-IRIS [144] introduced a hybrid Transformer that integrates tokens with stochastic $\Delta$-tokens to capture dynamics. $\text{D}^{2}$PO [145] employed preference learning to jointly optimize state prediction and action selection, enhancing the model’s understanding of underlying dynamics. For efficiency, MineWorld [59] accelerated token generation by predicting sequences in parallel and introduced an IDM as a controllability metric. Meanwhile, PIVOT-R [146] and ReOI [147] incorporated VLMs into control. PIVOT-R parses instructions to produce waypoint-based plans that an action module decodes into low-level controls, whereas ReOI detects implausible prediction elements, reimagines the distractors, and reintegrates the corrected content.
@@
Based on tokenization, some studies employ autoregressive diffusion to achieve stable generation and long-horizon planning. Epona [148] decouples spatiotemporal modeling, handled by a Transformer, from long-horizon multimodal generation, realized through trajectory and vision Diffusion Transformers (DiTs). Goff et al. [149] used a DiT to instantiate the state transition, which enables on-policy training and multi-second closed-loop rollouts. SceneDiffuser++[150] pushes this further to city-scale traffic simulation, applying multi-tensor diffusion over agents and traffic lights to produce stable closed-loop rollouts. For navigation, NWM [66] introduced an efficient conditional DiT to simulate visual trajectories for zero-shot planning.
@@
Another emerging direction is to inject explicit reasoning into the world model using LLMs and Chain-of-Thought (CoT). NavCoT [58] decomposes navigation into imagination, filtering, and prediction, enabling parameter-efficient in-domain training, and ECoT [54] leverages a pipeline of foundation models to generate reasoning labels for training a VLA policy. Variants like MineDreamer [79] introduced Chain-of-Imagination (CoI), where a multimodal LLM imagines future observations to steer diffusion and guide actions, and FSDrive [101] generates physics-constrained future scenes and treats them as CoT supervision, enabling VLMs to function as IDMs for planning.
@@
Other approaches directly couple LLMs with world models to operationalize planning and data generation. Dyna-Think [151] fuses reasoning and acting via a distilled LLM, and RIG [152] unifies reasoning and imagination end-to-end generalist policy. In terms of explicit dynamics and long-horizon, Gkountouras et al. [153] trained a causal world-model simulator that grounds an LLM nvironment causal reasoning and planning skills, Statler [46] enables LLMs to keep a structured world-state, using a reader for planning and a writer for updating, and Inner Monologue [42] incorporates a closed-loop feedback into LLMs, enabling agents to reason and deliberate more akin to human thinking. Finally, WoMAP [154] synthesized 3DGS scenes and trained a world model that refines VLM instructions for precise execution.
@@
In autonomous driving, many studies couple RNN-based dynamics with spatial grids to guide planning. For instance, DriveDreamer[91] and GenAD[92] adopt GRU-based dynamics over grid or instance-centric tokens to predict motion and decode trajectories. In contrast, DriveWorld [87] and Raw2Drive [155] instantiate RSSM dynamics on BEV tokens. DriveWorld conditions on tokens and actions for joint prediction, while Raw2Drive adopts a dual-stream design for spatiotemporal learning.
Numerous studies focus on autoregressively predicting future 3D occupancy representations to enable motion planning for autonomous driving. One strand discretizes scenes into occupancy tokens for sequential prediction, exemplified by OccWorld [93] and RenderWorld [156]. Another strand directly forecasts volumetric features or embeddings, as in Drive-OccWorld [157] and PreWorld [158]. Self-supervised variants predict future representations from current cues. For example, LAW [159] conditions on current representations and trajectories, SSR [160] compresses scenes into sparse BEV tokens for future BEV features, and NeMo [161] voxelizes multi-frame images and predicts occupancy to support imitation-based planning. Building on these representations, FASTopoWM [162] employs a unified decoder to align fast and slow systems from vehicle poses, enabling lane-topology reasoning, and WoTE [107] simulates candidate trajectories in BEV and selects among them with a reward model. Extending the paradigm, OccLLaMA [18] unifies occupancy, actions, and text within a single token vocabulary and employs a LLaMA for next token forecasting, planning, and question answering.
Beyond autonomous driving, similar formulations have been extended to broader domains of robotics. WMNav [163] leverages a VLM to maintain a curiosity-driven value map and adopts phased decision making to enable zero-shot, object-driven navigation. RoboOccWorld [164] targets indoor robotics by predicting fine-grained 3D occupancy with a pose-conditioned autoregressive Transformer, thereby supporting exploration and decision making. To achieve high-fidelity dynamics, EnerVerse [34] applies chunk-wise autoregressive video diffusion with a sparse memory mechanism to produce 4D latent dynamics and integrates 4DGS to mitigate the S2R gap in robotic execution. For manipulation, ParticleFormer [165] forecasts future point clouds with a Transformer-based particletized dynamics model, enabling robust handling of multi-object and multi-material interactions. At the representation level, DINO-WM [70] learns dynamics in the DINOv2 feature space and predicts future states to support zero-shot planning.
@@
Building on 3DGS, GAF [74] augments splats with learnable motion attributes to forecast future states and refines initial actions with diffusion. ManiGaussian [53] predicts per-point variations to generate future Gaussian scenes for manipulation under current states and actions, and ManiGaussian++ [80] adds a hierarchical leader-follower design with task-oriented splats to model primitive deformations for multi-body and bimanual skills. Within simulation and digital twin coupling, DreMa [60] integrates GS with a physics simulator to build twins for data synthesis in imitation learning, Abou-Chakra et al. [166] introduced a dual Gaussian–Particle representation where Gaussian points are attached to particles driven by visual loss forces, DexSim2Real2 [167] builds twins of articulated objects with generative models and uses sampling based planning for precise manipulation, PIN-WM [168] combines 3DGS with differentiable physics to estimate physical parameters from limited observations and generates digital cousins for zero-shot S2R policy learning, and PWTF [169] constructs an interactive twin that simulates candidate action outcomes and uses VLMs for evaluation and selection. At the representation level, DTT [98] adopts a triplane representation with multiscale Transformers to autoregressively capture incremental changes, forming a 4D world model for prediction and planning.
3.1.2 Global Difference Prediction

### v2

@@ III-A1 Sequential Simulation and Inference
Ha and Schmidhuber [ha2018recurrent] introduced an early world model that encodes observations into a latent space and employs an RNN to model dynamics for policy optimization. Building on this, PlaNet [hafner2019learning] introduced the Recurrent State-Space Model (RSSM), which fuses deterministic memory with stochastic components to enable robust long-horizon imagination. The successor models Dreamer, DreamerV2, and DreamerV3 [hafner2020dream, hafner2021mastering, hafner2025mastering] further advanced this formulation, inspiring a broad line of subsequent research.
@@
Building on RSSM, several variants modify or eliminate the decoder to better capture dynamics. For example, Dreaming [okada2021dreaming] uses contrastive learning and linear methods to mitigate state shifts, whereas DreamerPro [deng2022dreamerpro] replaces the decoder with prototypes to suppress visual distractions. To further enhance robustness, HRSSM [sun2024learning] was proposed, featuring a dual-branch architecture that aligns latent observations and shares information without reconstruction. Beyond architectural refinements, DisWM [wang2025disentangled] disentangles semantic knowledge from video content, distilling it into a world model that enables cross-domain generalization.
@@
A unifying theme across recent RSSM extensions is transferability, which reflects generalization across modalities, tasks, and embodiments for robust real-world robotics. At the representation level, PreLAR [zhang2024prelar] learns implicit action abstractions to bridge video-pretrained representations and control fine-tuning. Similarly, Wang et al. [wang2025latent] used optical flow as an embodiment-agnostic action representation to refine behavior-cloned policies, facilitating transfer across diverse embodiments. SENSEI [sancaktar2025sensei] distilled a Vision-Language Model (VLM) to derive semantic rewards and employed an RSSM that learns to predict and propagate these rewards internally. Under limited supervision, SR-AIF [nguyen2025sr] exploits prior preference learning and self-revision to enable adaptive learning in sparse-reward, continuous-control settings. To mitigate the Sim-to-Real (S2R) gap, ReDRAW [lanier2025adapting] is pretrained in simulation and adapted to the real environment using a limited amount of reward-free data, applying residual corrections to the latent dynamics. To handle mismatches, AdaWM [wang2025adawm] identifies discrepancies between the learned dynamics and the planner and selectively fine-tunes critical components. Other methods like WMP [lai2025world] address S2R transfer for challenging tasks, and DayDreamer [wu2023daydreamer] demonstrated sample-efficient deployment on physical robots. To broaden transfer, FOUNDER [wang2025founder] grounds representations from foundation models in the world-model state space, using temporal-distance prediction to handle flexible goals, and LUMOS [nematollahi2025lumos] introduced a language-conditioned imitation framework that operates on-policy in latent space with intrinsic rewards, enabling zero-shot transfer to real-world robotics.
@@
RSSM-based models have also been developed for autonomous driving. MILE [hu2022model] leverages offline expert data to enable imagined future states for planning. SEM2 [gao2024enhance] integrates semantic filtering with multi-source sampling to extract driving-relevant features and balance data distribution. Popov et al. [popov2025mitigating] addressed covariate shift through a latent generative world model that realigns policies with expert states. For safety, VL-SAFE [qu2025vl] supervises world models using safety scores derived from a VLM to generate safe trajectories. Finally, CALL [wang2025ego] extended the RSSM framework to Multi-Agent RL by introducing ego-centric information sharing to enhance planning capabilities.
@@
In contrast to RSSM, TransDreamer [chen2022transdreamer] introduced a Transformer State-Space Model (TSSM) that replaces the recurrent core in Dreamer, thereby substantially enhancing its capacity to capture long-horizon dependencies. The complementary OSVI-WM [goswami2025osvi] employs a causal Transformer for one-shot Imitation Learning (IL), autoregressively predicting the future latent trajectory and decoding it into physical waypoints for robot control.
@@
Some approaches continue to employ RNNs to capture temporal dependencies. On the modeling side, RWM [li2025robotic] introduced a dual-autoregressive, domain-agnostic neural simulator for long-horizon prediction. X-MOBILITY [liu2025x], in contrast, disentangles modeling from policy learning, using multi-head decoders for large-scale pretraining followed by supervised fine-tuning to derive strong policies. For humanoid locomotion, DWL [gu2024advancing] and WMR [sun2025learning] adopted end-to-end (E2E) frameworks. These frameworks reconstruct states from partial observations using either denoising or a gradient-blocked state estimator, which enables zero-shot transfer across complex real-world terrains.
@@
Recently, State Space Models (SSMs), exemplified by Mamba, have emerged as alternatives to RNNs and Transformers, combining linear-time complexity with long-horizon modeling capacity. Building on this, GLAM [he2025glam] improves both fidelity and efficiency via a Mamba-based parallel framework that integrates global and local modules to capture contextual and fine-grained dynamics.
@@
Beyond forward temporal modeling, Inverse Dynamics Modeling (IDM) is a key paradigm in world model construction. An IDM infers the actions required to transition between an initial and a target state. Agrawal et al. [agrawal2016learning] integrated a forward model with an IDM for multi-step prediction, establishing the basis for subsequent research. More recent work includes GLAMOR [paster2021planning], which trains an object-conditioned IDM to predict the actions necessary to reach a specified target. In Dreamer-style agents, Iso-Dream[pan2022iso] leverages IDM to decompose the world model into controllable and uncontrollable components, using the rollout of uncontrollable states to guide policy learning.
@@
Recent RSSM-centric studies have begun to exploit token-level dependencies to strengthen representation learning and temporal reasoning. For example, MWM [seo2023masked] decouples visual tokens from RSSM-based dynamics via a masked autoencoder, improving both performance and data efficiency. NavMorph [yao2025navmorph] introduced a self-evolving RSSM with a contextual evolution memory for online adaptation. For temporal abstraction, WISTER [burchi2025learning] employed action-conditioned contrastive predictive coding to train a TSSM that captures high-level temporal features. Similarly, TWM [robine2023transformerbased] used a Transformer to align multimodal tokens with historical states during training, while relying on a lightweight policy at inference. To handle long-horizon tasks, some approaches integrate LLMs with RSSMs to decompose objectives into subtasks. EvoAgent [feng2025evoagent], for example, uses an LLM to guide low-level actions and regularizes RSSM updates. RoboHorizon [chen2025robohorizon], in contrast, enhances task recognition with dense rewards and leverages key task segments via a masked autoencoder.
@@
In autonomous driving, token-based sequence representations are increasingly adopted to model cross-modal interactions and spatiotemporal structures. DrivingWorld [hu2024drivingworld] pairs next-state prediction for temporal dynamics with next-token prediction for spatial structure. For multimodal control, Doe-1 [zheng2024doe] formulates closed-loop driving as autoregressive prediction over perception-description-action tokens, which unifies perception, prediction, and planning, and DrivingGPT [chen2024drivinggpt] interleaves vision and action tokens and casts world modeling and trajectory planning as next-token prediction. To enhance diversity and safety, LatentDriver [xiao2025learning] models future actions as a mixture distribution and actuates the world model with planner-sampled intermediate actions. At the same time, Vasudevan et al. [vasudevan2025planning] proposed an adaptive model that predicts surrounding agents for safe navigation.
@@
The token-based paradigm also extends to broader robotics. Within RL, IRIS [micheli2023transformers] and TWM [dedieu2025improving] leverage discrete tokens to enable data-efficient policy learning via imagined or hybrid rollouts. DyWA [lyu2025dywa] improves action learning by conditioning on trajectory dynamics and jointly predicting future states with single-view point-cloud and proprioceptive modalities. EgoAgent [chen2025egoagent] interleaves state-action sequence modeling within a Transformer, enabling unified perception, prediction, and action inference. Tokenized representations unify multimodal inputs, including vision, language, and action (VLA), enabling generalist agents with cross-domain adaptability, as shown in WorldVLA [cen2025worldvla]. Recent studies encode environmental states as discrete symbolic tokens and condition next-token prediction on action, as demonstrated by DCWM [scannell2025discrete] and TrajWorld [yin2025trajectory].
@@
Recent studies have strengthened the link between tokenized representations and planning, particularly through object-centric approaches. These models, such as CarFormer [hamdan2024carformer], the work of Jeong et al. [jeong2025object], and Dyn-O [wang2025dyn], represent scenes as a collection of slots. CarFormer autoregressively models the relationships between these slots in BEV. Jeong et al. added language-guided manipulation, while Dyn-O used a Mamba with dropout scheduling for temporal modeling and to disentangle static from dynamic elements. $\Delta$-IRIS [micheli2024efficient] introduced a hybrid Transformer that integrates tokens with stochastic $\Delta$-tokens to capture dynamics. $\text{D}^{2}$PO [wang2025world] employed preference learning to jointly optimize state prediction and action selection, enhancing the model’s understanding of underlying dynamics. For efficiency, MineWorld [guo2025mineworld] accelerated token generation by predicting sequences in parallel and introduced an IDM as a controllability metric. Meanwhile, PIVOT-R [zhang2024pivot] and ReOI [chen2025reimagination] incorporated VLMs into control. PIVOT-R parses instructions to produce waypoint-based plans that an action module decodes into low-level controls, whereas ReOI detects implausible prediction elements, reimagines the distractors, and reintegrates the corrected content.
@@
Based on tokenization, some studies employ autoregressive diffusion to achieve stable generation and long-horizon planning. Epona [zhang2025epona] decouples spatiotemporal modeling, handled by a Transformer, from long-horizon multimodal generation, realized through trajectory and vision Diffusion Transformers (DiTs). Goff et al. [goff2025learning] used a DiT to instantiate the state transition, which enables on-policy training and multi-second closed-loop rollouts. SceneDiffuser++[tan2025scenediffuser++] pushes this further to city-scale traffic simulation, applying multi-tensor diffusion over agents and traffic lights to produce stable closed-loop rollouts. For navigation, NWM [bar2025navigation] introduced an efficient conditional DiT to simulate visual trajectories for zero-shot planning.
@@
Another emerging direction is to inject explicit reasoning into the world model using LLMs and Chain-of-Thought (CoT). NavCoT [lin2025navcot] decomposes navigation into imagination, filtering, and prediction, enabling parameter-efficient in-domain training, and ECoT [zawalski2025robotic] leverages a pipeline of foundation models to generate reasoning labels for training a VLA policy. Variants like MineDreamer [zhou2025minedreamer] introduced Chain-of-Imagination (CoI), where a multimodal LLM imagines future observations to steer diffusion and guide actions, and FSDrive [zeng2025futuresightdrive] generates physics-constrained future scenes and treats them as CoT supervision, enabling VLMs to function as IDMs for planning.
@@
Other approaches directly couple LLMs with world models to operationalize planning and data generation. Dyna-Think [yu2025dyna] fuses reasoning and acting via a distilled LLM, and RIG [zhao2025rig] unifies reasoning and imagination end-to-end generalist policy. In terms of explicit dynamics and long-horizon, Gkountouras et al. [gkountouras2025language] trained a causal world-model simulator that grounds an LLM nvironment causal reasoning and planning skills, Statler [yoneda2024statler] enables LLMs to keep a structured world-state, using a reader for planning and a writer for updating, and Inner Monologue [huang2023inner] incorporates a closed-loop feedback into LLMs, enabling agents to reason and deliberate more akin to human thinking. Finally, WoMAP [yin2025womap] synthesized 3DGS scenes and trained a world model that refines VLM instructions for precise execution.
@@
In autonomous driving, many studies couple RNN-based dynamics with spatial grids to guide planning. For instance, DriveDreamer[wang2024drivedreamer] and GenAD[zheng2024genad] adopt GRU-based dynamics over grid or instance-centric tokens to predict motion and decode trajectories. In contrast, DriveWorld [min2024driveworld] and Raw2Drive [yang2025raw2drive] instantiate RSSM dynamics on BEV tokens. DriveWorld conditions on tokens and actions for joint prediction, while Raw2Drive adopts a dual-stream design for spatiotemporal learning.
Numerous studies focus on autoregressively predicting future 3D occupancy representations to enable motion planning for autonomous driving. One strand discretizes scenes into occupancy tokens for sequential prediction, exemplified by OccWorld [zheng2024occworld] and RenderWorld [yan2025renderworld]. Another strand directly forecasts volumetric features or embeddings, as in Drive-OccWorld [yang2025driving] and PreWorld [li2025semi]. Self-supervised variants predict future representations from current cues. For example, LAW [li2025enhancing] conditions on current representations and trajectories, SSR [li2025navigation] compresses scenes into sparse BEV tokens for future BEV features, and NeMo [huang2024neural] voxelizes multi-frame images and predicts occupancy to support imitation-based planning. Building on these representations, FASTopoWM [yang2025fastopowm] employs a unified decoder to align fast and slow systems from vehicle poses, enabling lane-topology reasoning, and WoTE [li2025end] simulates candidate trajectories in BEV and selects among them with a reward model. Extending the paradigm, OccLLaMA [wei2024occllama] unifies occupancy, actions, and text within a single token vocabulary and employs a LLaMA for next token forecasting, planning, and question answering.
Beyond autonomous driving, similar formulations have been extended to broader domains of robotics. WMNav [nie2025wmnav] leverages a VLM to maintain a curiosity-driven value map and adopts phased decision making to enable zero-shot, object-driven navigation. RoboOccWorld [zhang2025occupancy] targets indoor robotics by predicting fine-grained 3D occupancy with a pose-conditioned autoregressive Transformer, thereby supporting exploration and decision making. To achieve high-fidelity dynamics, EnerVerse [huang2025enerverse] applies chunk-wise autoregressive video diffusion with a sparse memory mechanism to produce 4D latent dynamics and integrates 4DGS to mitigate the S2R gap in robotic execution. For manipulation, ParticleFormer [huang2025particleformer] forecasts future point clouds with a Transformer-based particletized dynamics model, enabling robust handling of multi-object and multi-material interactions. At the representation level, DINO-WM [zhou2025dino] learns dynamics in the DINOv2 feature space and predicts future states to support zero-shot planning.
@@
Building on 3DGS, GAF [chai2025gaf] augments splats with learnable motion attributes to forecast future states and refines initial actions with diffusion. ManiGaussian [lu2024manigaussian] predicts per-point variations to generate future Gaussian scenes for manipulation under current states and actions, and ManiGaussian++ [yu2025manigaussian++] adds a hierarchical leader-follower design with task-oriented splats to model primitive deformations for multi-body and bimanual skills. Within simulation and digital twin coupling, DreMa [barcellona2025dream] integrates GS with a physics simulator to build twins for data synthesis in imitation learning, Abou-Chakra et al. [abou2025physically] introduced a dual Gaussian–Particle representation where Gaussian points are attached to particles driven by visual loss forces, DexSim2Real2 [jiang2025dexsim2real] builds twins of articulated objects with generative models and uses sampling based planning for precise manipulation, PIN-WM [li2025pin] combines 3DGS with differentiable physics to estimate physical parameters from limited observations and generates digital cousins for zero-shot S2R policy learning, and PWTF [ning2025prompting] constructs an interactive twin that simulates candidate action outcomes and uses VLMs for evaluation and selection. At the representation level, DTT [xu2025delta] adopts a triplane representation with multiscale Transformers to autoregressively capture incremental changes, forming a 4D world model for prediction and planning.
III-A2 Global Difference Prediction